THE PURE ROTATIONAL SPECTRUM OF CuCH_{3}

\underline{D. T. HALFEN}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721; D. B. GROTJAHN, Department of Chemistry, San Diego State University, San Diego, CA, 92182; and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of CuCH_{3} (X̃^{1}A_{1}) has been measured with millimeter-wave direct absorption techniques in the frequency range of 312 to 543 GHz. This work is the first spectroscopic observation of monomeric CuCH_{3} in the gas-phase. Measurements have been made of the two Cu isotopomers, ^{63}Cu and ^{65}Cu, in the ground and v_{3}=1 (the Cu-C stretch) vibrational states. Copper monomethyl was formed by the reaction of Cu vapor produced in a Broida-type oven and Sn(CH_{3})_{4}. Ten transitions of both isotopomers have been recorded with K components up to 13. Additional work is being conducted on Cu^{13}CH_{3} and CuCD_{3}. From this data, spectroscopic constants and structural information are being determined. Establishing the geometry of copper monomethyl is significant for organometallic chemistry.